OpenStudy (anonymous):

y=-2x2 + 28x - 26. Factored form

hero (hero):

y=-2x2 + 28x - 26 To properly factor this, set y = 0 and solve for x as follows: -2x^2 + 28x - 26 = 0 2x^2 - 28x + 26 = 0 2(x^2 - 14x + 13) = 0 x^2 - 14x + 13 = 0 x^2 - 13x - 1x + 13 = 0 x(x-13)-1(x-13) = 0 (x-13)(x-1) = 0